OpenStudy (anonymous):

Jack usually mows his lawn in 6 hours. Marilyn can mow the same yard in 4 hours. How much time would it take for them to mow the lawn together?

OpenStudy (anonymous):

(1/6+1/4=1/x )= 4x+6x=24 = change the denominators to a common, and just solve the equation.

OpenStudy (anonymous):

Can't u just tell me the answer?

OpenStudy (zarkon):

just solve 4x+6x=24 for x

OpenStudy (turingtest):

We CAN just tell you the answer, but then we feel like you're not learning anything. You wouldn't want that now, would you?

OpenStudy (anonymous):

Yeah actually i would. :D

OpenStudy (turingtest):

Well, if you wait around you'll eventually find those tutors who just want medals, so you'll have to hold out for that. I suggest you at least try to solve 4x+6x=24 Please show us your work, and if you get it wrong we usually have no problem explaining where you made your mistake.

OpenStudy (anonymous):

is it 12/5

OpenStudy (turingtest):

bingo! not that hard, eh? have a medal!

OpenStudy (anonymous):

not with Mathway.com. ;)

OpenStudy (turingtest):

That's cute. If you want to know the grown-up way to cheat like that use WolframAlpha, It can do all kinds of that stuff. I do wish I could revoke your medal though :/

OpenStudy (turingtest):

http://www.wolframalpha.com/input/?i=4x%2B6x%3D24

OpenStudy (anonymous):

lol Thx 4 the tip.